If 0.200 moles of AgNO3 react with 0.155 moles of H2SO4 according to this UNBALANCED equation below, how many grams of AgSO4 could be formed? AgNO3+H2SO4->Ag2SO4+HNO3
1
Expert's answer
2020-03-25T12:48:36-0400

The balanced chemical equation can be written as follows:

2AgNO3+H2SO4->Ag2SO4+2HNO3

Since 2 moles of AgNO3 react with 1 mole of H2SO4 to form 1 mole of Ag2SO4, n (Ag2SO4):


"n=0.200\times\frac{0.155}{2}\times\frac{1}{1}=0.0155"


"m=M\times n=0.0155\times311.7984=4.83"